OpenStudy (anonymous):

can someone tell me if i am correct? sqrt(x+8)=-4 i got x=8

OpenStudy (lukecrayonz):

To remove the sqrt, you square both sides. x+8=16, as -4*-4 = 16. x+8=16 -8 both sides x=8

OpenStudy (anonymous):

that's what i thought, but the system is telling me i am wrong :/

OpenStudy (lukecrayonz):

Ohh, I see. If you replace 8 as x, then sqrt(16)=-4, and sqrt16= positive 4.

OpenStudy (lukecrayonz):

It's the sign change from the squaring.

OpenStudy (anonymous):

lol, i think i found a system mistake, says the answer is B rofl

OpenStudy (anonymous):

b is no solution